Hall C meeting 1/4/10- 3p.m. Qweak: Greg Smith Mapper picked up

Hall C meeting
1/4/10- 3p.m.
Qweak: Greg Smith
Mapper picked up from Bates and is en route, will be delivered directly to Hall.
Quartz scanner arrived before Christmas and is in storage in EEL 126
Milestone meeting: Next week
Collaboration meeting: Jan 31-Feb2
MCC coordination meeting: Tues at 8:30
Onsite Qweak meeting- 1/15 at 1:30 pm
Work Coordinator: Kenyon
2nd coil installed but awaiting alignment (alignment group busy with injector/arc 10 work), 1/3 will be
installed post alignment.
Laying of concrete forms this week
A design is required of 4 more feet for the ferris wheel (within this month)
Work on SOS post Qweak (plan it, SOS free to move)
-AC to SOS disconnected today
-test cut on concrete (wet cutting) – cost estimate/proposal by mid jan.
-Estimated removal cost- 500 k, ~50K for dipole, ~120k for frame
Engineering: Joe Brindza
No real change in power supply since shut down, mapping forthcoming.
For counting house: New 50” monitor and rebuilt batteries for UPS’s (currently on line)
NAND power supplies with cost 50% of their new cost to rebuild
MIT power bus was found to have been left outside, and as such, corroding over down; custom cabling
for Bus supply needs to be located.
Target status:
Jim Dunn on motion, cleaning up relay boxes (very buggy)
Slow progress is being made.
Help is required for vent line installation as designed by Gary Vasilaro, a suggestion was made to see if
this falls under the realm of the Lockwood general labor contract. The loop is being fabricated and flex
hoses are en route. A workable design for the insulating window seal is as yet incomplete.
Shield connections for Moeller magnet and vent line in progress, for 12 GeV a boron carbide neutron
sink has been produced which is packing to 90% of theoretical efficiency with only the use of hand tools.
Combined procurement proposal has been canceled and is breaking down into 2 proposals; the magnet
and everything else. On the Bender magnet, Michigan state has an offer on the table which is currently
under discussion, but will likely be proceeded with.
The Q1 magnet will need to be sent for bid again (3 months) due to contractor pull out. This will cost
some of the float time in the proposal. It should hopefully be confirmed by late March (before the April
Lehmann review?)
Better chain/hand rails need be installed on the mezzanine for safety reasons
PAC35 meeting:
4 TAC proposals and a letter of intent.
The letter of intent was not present. TAC proposals reviewed were for PR10-003, PR10-002, PR10-12,
and PR10-008
Basic notes:
PR10-003: 21 days @ 11 GeV and 80 uA. HMS is the proton arm in the interaction. Contact Werner
Boeglin. “Deuteron electro-disintegration …”
PR-10-002: 13 days @ 6.6/11 GeV at 40 or less uA. “Inclusive scattering: Precise measurement of the F2
structure function at large x in the resonance region and beyond”
PR10-12: “Precision measurement of nucleon and nuclear structure functions to constrain gluon
distribution”, 11 days at 4 different energies at 40uA each (one non-standard 7.7 GeV). There was
discussion of the addition of an iron target as well as a review of the systematic errors of a similar
previous experiment.
PR10-008: “Detailed studies of the nuclear dependence of F2 in light nuclei” 23 days at 11 GeV at 15-80
uA. Note of the expense and difficulties associated with the Ca-48 target as well as the Li-6 target were